OpenStudy (anonymous):

A sail is in the form of a right triangle that is four times as high as it is wide. The sail is made from 32 square meters of material.

OpenStudy (imstuck):

im guessing you have to find the height andd the width of the sail?

OpenStudy (anonymous):

just the height

OpenStudy (imstuck):

Ok, so if the height is 4 times the width, it takes 4 widths to equal 1 height, so h = 4w. We will use w instead of b for the base, ok?

OpenStudy (imstuck):

The formula for the area of a triangle is A = 1/2 * w * h. We know the area is 32, and we have a direct relationship between the width and the height. So we can substitute in what we know now. Like this:

OpenStudy (anonymous):

yeah

OpenStudy (imstuck):

\[32=\frac{ 1 }{ 2 }(w)(4w)\]

OpenStudy (imstuck):

We can now solve for w. Then from that we can find the height.

OpenStudy (imstuck):

multiply both sides by 2 to get rid of the fraction.

OpenStudy (anonymous):

only find the height

OpenStudy (imstuck):

I know we only need the height, but you can't find the height without finding the width. So bear with me here.

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

When you multiply both sides by 2, you get this:\[64=4w ^{2}\]

OpenStudy (imstuck):

Divide both sides by 4 to get \[16=w ^{2}\]now what is w?

OpenStudy (anonymous):

4

OpenStudy (imstuck):

Good! Now remember that we said the height was 4 times the width. Well the width is 4, so what is the height? h = 4w, so h = 4(4).